# Quantum entanglement between partons in a strongly coupled quantum field theory

### v1

@@ 3 Light-front wave function representation
In the quenched approximation, the valence, normalized, and one-parton PDFs all coincide with $f^{\text{Que}}_{N}(x)$. Fig. 2 compares these distributions in both quenched and unquenched calculations. At moderate coupling $\alpha=1.0$, the differences are negligible; at strong coupling $\alpha=2.0$, deviations become visible, reflecting the contribution of sea partons.
@@ 4 Density matrices
Periodic boundary conditions (PBC) are imposed in the transverse directions, while in the longitudinal direction we apply PBC for bosons and anti-periodic boundary conditions (ABC) for fermions. The allowed momenta become discrete:
@@
It is convenient to introduce normalized momentum states
$|p)=(2p^{+}V)^{-\frac{1}{2}}|p\rangle,$ (48)
which satisfy $(p^{\prime}|p)=\delta_{p,p^{\prime}}$. The continuum limit is recovered as $L,L_{\perp}\to\infty$. For a hadron with definite longitudinal momentum $P_{0}^{+}$, we set $P_{0}^{+}=(2\pi/L)K$, where $K$ is known as the light-cone harmonic resolution Pauli:1985ps ; Pauli:1985pv . Light-front boost invariance ensures that the intrinsic LFWFs depend only on the longitudinal momentum fractions $x_{i}=n_{i}/K$ and relative momentum $\vec{k}_{i\perp}$. In the continuum limit, $K\to\infty$ while $K/L$ remains finite. In what follows, we perform calculations in the discretized framework and take the continuum limit at the end.
@@ 5 Entanglement in quenched scalar theory
Taking into account the IR divergence from the one-body term and the fact that $f_{N}^{(n)}(x\to 1,{k}_{\perp})\to 0$ for $n\geq 2$, the entanglement entropy becomes
@@
The corresponding discrete form, for finite volume $V$, is
@@ 6 Entanglement in unquenched scalar theory
$S_{N}\equiv S_{\text{vN}}(\rho_{N})=S_{\text{vN}}(\rho_{N}^{(1)})+S_{\text{vN}}(\rho_{N}^{(2)}).$ (74)
@@
$S_{\text{vN}}(\varrho_{N})=S_{\text{vN}}(\varrho_{\bar{N}}),$ (75)
@@
$S_{\text{vN}}(\varrho_{\bar{N}})=Z_{NN\bar{N}}\log(P_{0}^{+}V)-\int{\mathrm{d}}x\int\frac{{\mathrm{d}}^{2}k_{\perp}}{(2\pi)^{3}}f_{\bar{N}}(x,{k}_{\perp})\log f_{\bar{N}}(x,{k}_{\perp}).$ (78)
@@ 7 Discussions
From an experimental standpoint, the density matrix, and hence the entanglement entropy, is only accessible up to a finite factorization scale $\mu_{F}$. Modes with transverse momentum $k_{\perp}\gtrsim\mu_{F}$ are unresolved and must be traced out. This motivates the definition of a scale-dependent reduced density matrix:
@@
While quantum negativity can be computed numerically in our scalar Yukawa model, we can gain deeper physical insight by considering the physical nucleon within a two-body truncation. In this limit, the PPT density matrix can be diagonalized analytically, yielding the following eigenvalues:
$Z,\;\pm\sqrt{\frac{Z}{P_{0}^{+}V}}\frac{\psi_{\pi N}(x,k_{\perp})}{\sqrt{2x(1-x)}},\;\frac{1}{P_{0}^{+}V}\frac{|\psi_{\pi N}(x,k_{\perp})|^{2}}{2x(1-x)},\;\pm\frac{1}{P_{0}^{+}V}\frac{\psi^{*}_{\pi N}(x^{{}^{\prime}},k^{{}^{\prime}}_{\perp})\psi_{\pi N}(x,k_{\perp})}{\sqrt{2x(1-x)}\sqrt{2x^{{}^{\prime}}(1-x^{{}^{\prime}})}}.$ (103)
The presence of negative eigenvalues confirms the existence of entanglement, consistent with our previous entropy results. [^2] We note that within the two-body truncation, the $N\pi$ system is a pure state, for which entanglement entropy is also a valid measure. Furthermore, the corresponding quantum negativity takes a compact form:
@@ Acknowledgements
This work was supported in part by the Chinese Academy of Sciences under Grant No. YSBR-101, by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081. WQ is supported by the European Research Council under project ERC-2018-ADG-835105 YoctoLHC; by Maria de Maeztu excellence unit grant CEX2023-001318-M and project PID2020-119632GB-I00 funded by MICIU/AEI/10.13039/501100011033; by ERDF/EU; by the Marie Sklodowska-Curie Actions Fellowships under Grant No. 101109293; and by Xunta de Galicia (CIGUS Network of Research Centres). Y.Z. is supported by the European Union “Next Generation EU” program through the Italian PRIN 2022 grant n. 20225ZHA7W.

### v2

@@ 3 Light-front wave function representation
In the quenched approximation, the valence, normalized, and one-parton PDFs all coincide with $f^{\text{Que}}_{N}(x)$. Fig. 2 compares these distributions in both quenched and unquenched calculations. At moderate coupling $\alpha=1.0$, the differences are negligible; at strong coupling $\alpha=2.0$, deviations become visible, reflecting the contribution of sea partons. As we will show later in Sec. 6 that the Shannon entropy of $\hat{f}(x,k_{\perp})$ and $\bar{f}(x,k_{\perp})$ are different from the parton entanglement entropy.
@@ 4 Density matrices
We follow the standard convention in discretized light-cone quantization (DLCQ) and adopt the periodic boundary conditions (PBC) in the transverse directions, while in the longitudinal direction we apply PBC for bosons and anti-periodic boundary conditions (ABC) for fermions Brodsky:1997de . The allowed momenta become discrete:
@@
The continuum limit is recovered as $L,L_{\perp}\to\infty$. For a hadron with definite longitudinal momentum $P_{0}^{+}$, we set $P_{0}^{+}=(2\pi/L)K$, where $K$ is known as the light-cone harmonic resolution Pauli:1985ps ; Pauli:1985pv . Light-front boost invariance ensures that the intrinsic LFWFs depend only on the longitudinal momentum fractions $x_{i}=n_{i}/K$ and relative momentum $\vec{k}_{i\perp}$. In the continuum limit, $K\to\infty$ while $K/L$ remains finite. In what follows, we perform calculations in the discretized framework and take the continuum limit at the end.
@@ 5 Entanglement in quenched scalar theory
Substituting (59) into (56), we obtain
$S_{N}=\log(P_{0}^{+}V)-\int_{0}^{1}\text{d}x\int\frac{\text{d}^{2}k_{\perp}}{(2\pi)^{3}}\Big[f_{N}^{(1)}(x,k_{\perp})+\sum_{n\geq 2}f_{N}^{(n)}(x,k_{\perp})\Big]\\ \times\log\Big[f_{N}^{(1)}(x,k_{\perp})+\sum_{n\geq 2}f_{N}^{(n)}(x,k_{\perp})\Big].$ (61)
Taking into account the fact that $f_{N}^{(1)}(x,{k}_{\perp})=Z(2\pi)^{3}\delta(x-1)\delta^{2}({k}_{\perp})$ and $f_{N}^{(n)}(x\to 1,{k}_{\perp})\to 0$ for $n\geq 2$, the $n=1$ term can be analytically evaluated and the entanglement entropy becomes,
@@
Here, the $f_{N}^{(1)}(x,k_{\perp})$ in the logarithm in (61) needs to use the discretized version $f_{N}^{(1)}(x,k_{\perp})$ $ZP_{0}^{+}V\delta_{p,P_{0}}$. The corresponding discrete form for finite volume $V$ is
@@ 6 Entanglement in unquenched scalar theory
$S_{N}\equiv S_{\text{vN}}(\rho_{N})=-\operatorname{Tr}(\rho_{N}^{(1)}\log\rho_{N}^{(1)})-\operatorname{Tr}(\rho_{N}^{(2)}\log\rho_{N}^{(2)}).$ (74)
@@
$\operatorname{Tr}(\varrho_{N}\log\varrho_{N})=\operatorname{Tr}(\varrho_{\bar{N}}\log\varrho_{\bar{N}}),$ (75)
@@
$\operatorname{Tr}(\varrho_{\bar{N}}\log\varrho_{\bar{N}})=-Z_{NN\bar{N}}\log(P_{0}^{+}V)+\int{\mathrm{d}}x\int\frac{{\mathrm{d}}^{2}k_{\perp}}{(2\pi)^{3}}f_{\bar{N}}(x,{k}_{\perp})\log f_{\bar{N}}(x,{k}_{\perp}).$ (78)
@@ 7 Discussions
In experiments, the transverse momentum of a parton is only observable within the finite acceptance of the detector. Therefore, the density matrix, and hence the entanglement entropy, is only accessible up to a finite factorization scale $\mu_{F}$. Modes with transverse momentum $k_{\perp}\gtrsim\mu_{F}$ are unresolved and must be traced out. This motivates the definition of a scale-dependent reduced density matrix:
@@
While quantum negativity can be computed numerically in our scalar Yukawa model, we can gain deeper physical insight by considering the physical nucleon within a two-body truncation $|N\rangle+|N\pi\rangle$, in which the Fock space can be bipartitioned into the nucleons and the pions (including zero-pions). In this limit, the PPT density matrix can be diagonalized analytically. In the quenched case, the partially transposed density matrix $\rho^{T_{N}}$ can be obtained by partially transposing the Hilbert space $\mathcal{H}_{N}$ of the density matrix
$\begin{split}\rho^{T_{N}}=\sum_{{\mathcal{F}},{\mathcal{G}}}&\frac{1}{S_{\mathcal{F}}S_{\mathcal{G}}}\prod_{i\in{\mathcal{F}}}\int\frac{{\mathrm{d}}^{3}p^{\prime}_{i}}{(2\pi)^{3}2p^{\prime+}_{i}}\int\frac{{\mathrm{d}}^{3}p^{\prime}_{N}}{(2\pi)^{3}2p^{\prime+}_{N}}\prod_{j\in{\mathcal{G}}}\int\frac{{\mathrm{d}}^{3}p_{j}}{(2\pi)^{3}2p_{j}^{+}}\\ &\int\frac{{\mathrm{d}}^{3}p_{N}}{(2\pi)^{3}2p_{N}^{+}}\Psi_{\mathcal{F}}^{*}(\{p^{\prime}_{i}\}_{\mathcal{F}},p_{N}^{\prime})\Psi_{\mathcal{G}}(\{p_{j}\}_{\mathcal{G}},p_{N})|\{p^{\prime}_{i}\}_{\mathcal{F}},p_{N}\rangle\langle\{p_{j}\}_{\mathcal{G}},p_{N}^{\prime}|.\end{split}$ (103)
Here $\mathcal{F}$ and $\mathcal{G}$ only contain pions. In the two-body truncation, by adopting (42), (46) and (47) to $\rho^{T_{N}}$, the partially transposed density matrix can be written as
[…]
To solve the eigenvalue equation,
$\rho^{T_{N}}_{2-\text{body}}|\phi\rangle=\lambda|\phi\rangle,$ (105)
we first discretize the momenta and project the equation to one- and two-body Fock sectors, which yields the following set of equations,
$\begin{gathered}\lambda_{1}=Z\\ \lambda_{2}^{\pm}(x,k_{\perp})=\pm\sqrt{\frac{Z}{P_{0}^{+}V}}\frac{|\psi_{\pi N}(x,k_{\perp})|}{\sqrt{2x(1-x)}}\\ \lambda_{3}(x,k_{\perp})=\frac{1}{P_{0}^{+}V}\frac{|\psi_{\pi N}(x,k_{\perp})|^{2}}{2x(1-x)}\\ \lambda_{4}^{\pm}(x,k_{\perp},x^{\prime},k_{\perp}^{\prime})=\pm\frac{1}{P_{0}^{+}V}\frac{|\psi_{\pi N}(x^{\prime},k^{\prime}_{\perp})\psi_{\pi N}(x,k_{\perp})|}{\sqrt{2x(1-x)}\sqrt{2x^{\prime}(1-x^{\prime})}}\end{gathered}$ (106)
for $x,x^{\prime}\in(0,1)$ and $k_{\perp},k_{\perp}^{\prime}\in\mathbb{R}^{2}$. The presence of negative eigenvalues confirms the existence of entanglement, consistent with our previous entropy results. [^2] We note that within the two-body truncation, the $N\pi$ system is a pure state, for which entanglement entropy is also a valid measure. Furthermore, the corresponding quantum negativity takes a compact form:
@@ Acknowledgements
This work was supported in part by the Chinese Academy of Sciences under Grant No. YSBR-101, by the National Natural Science Foundation of China (NSFC) under Grant Nos. 12375081, 125B2111. WQ is supported by the European Research Council under project ERC-2018-ADG-835105 YoctoLHC; by Maria de Maeztu excellence unit grant CEX2023-001318-M and project PID2020-119632GB-I00 funded by MICIU/AEI/10.13039/501100011033; by ERDF/EU; by the Marie Sklodowska-Curie Actions Fellowships under Grant No. 101109293; and by Xunta de Galicia (CIGUS Network of Research Centres). Y.Z. is supported by the European Union “Next Generation EU” program through the Italian PRIN 2022 grant n. 20225ZHA7W.